OpenStudy (anonymous):

Having trouble solving this Laplace transform: y''−8y'+16y=t, y(0)=0, y'(0)=1 Here's what I have so far: y(s)(s-4)^2=1/s+1 y(s)=1/[(s^2)(s-4)^2]+1/(s-4)^2 partial decomp of 1/[(s^2)(s-4)^2]: 1/16(1/s^2)-1/8[1/(s-4)]+1/16[1/(s-4)^2] y(s)=1/16(1/s^2)-1/8[1/(s-4)]+17/16[1/(s-4)^2] y(t)=t/16-[e^(4t)]/8+[17te^(4t)]/16 But I get the wrong answer. Help?

OpenStudy (kainui):

Give me some time to refresh and work through this. It's been a while for me but I think I can help you out.

OpenStudy (anonymous):

Alright, thank you.

OpenStudy (kainui):

I think I found your problem. When you did partial fractions you needed to add an extra term: \[\Large \frac{1}{s^2(s-4)^2}=\frac{A}{s}+\frac{B}{s^2}+\frac{C}{s-4}+\frac{D}{(s-4)^2}\]

OpenStudy (anonymous):

Yeah, I did that. I got A=0. Maybe I did the partial fraction decomp wrong?

OpenStudy (kainui):

Yeah I think so, try it again. I think that's it.

OpenStudy (anonymous):

Alright, give me 2 min.

OpenStudy (kainui):

No problem, there's a lot going on in that mess haha.

OpenStudy (anonymous):

Still getting A=0.

OpenStudy (anonymous):

And yeah, there's way too much arithmetic in differential equations.

OpenStudy (kainui):

Hmm, I'm getting A=1/32. Ok, I guess it won't be very pleasant for me to type out how I got there.

OpenStudy (kainui):

How do you solve partial fractions?

OpenStudy (kainui):

I basically distribute everything out, collect terms of the same powers of s to create 4 equations. Then you can either put it in a matrix to solve it or whatever.

OpenStudy (anonymous):

Ahh, wait a minute. I put in Cs(s-4) instead of Cs^2(s-4). Let me rework it.

OpenStudy (kainui):

Cool =)

OpenStudy (anonymous):

You get A=1/32, B=1/16,C=-1/32,D=1/16?

OpenStudy (anonymous):

Just submitted answer. It's right!! Second time in two days partial fraction decomp has tripped me up. Anyways, thanks for the help. I really appreciate it.

OpenStudy (kainui):

Awesome, glad I could help. It wasn't even really the transform part that's hard haha.

OpenStudy (anonymous):

Yeah, I feel like this class tests me more on attention to detail than anything else.

OpenStudy (kainui):

Yeah, in a way these classes sort of battle hardened me to calculating things faster and more precisely, which is kind of good in a way. It's like you have to go through calculus to master algebra haha.